# Global Clipper: Enhancing Safety and Reliability of Transformer-based Object Detection Models

### v1

@@ 1 Motivation
Errors during inference can lead to faulty bounding boxes, significantly altering system behaviour and underscoring the critical need for safer hardware for model execution. DNN accelerators execute models at a high level by constructing a computational graph that uses General matrix-to-matrix multiplication (GEMM) [6] for extensive layer input and weight multiplications. Key components in this process are the Multiply-accumulate (MAC) units within the lower accelerator levels shown in fig. 1 [7]. MAC units in DNN accelerators lack ECC protection, making them particularly vulnerable to soft errors—a major reliability concern. Such errors, often caused by radiation, chip ageing, manufacturing variations, or thermal issues [8, 9, 10], can alter intermediary computational values, leading to incorrect inferences. Research shows that the soft error rate will increase with higher transistor density, reduced feature sizes, and more cores [11, 12, 13]. For example, a 100-core system may fail every 1.5 hours due to soft errors, significantly affecting predictions as these propagate through layers, as shown in fig. 1. Although soft errors do not cause permanent damage, they can result in substantial reliability degradation.
@@ 2 Related Work
The reliability of safety-critical DNN models is assessed through various metrics at the application and hardware levels, enhancing their safety and reliability [19, 20, 15, 18]. Research has shown that DNNs are prone to soft errors, with instances of single-bit flips leading to faulty inferences [21, 22, 23, 18]. Traditionally, protection from soft errors in hardware has primarily involved error detection or correction codes (EDC or ECC) [24] for memory and using residuals for computing. These mechanisms are commonly implemented in high-end server-grade CPUs but less so in GPUs due to cost considerations or typical relaxed application requirements. Other techniques based on redundancy, like DMR (dual modular redundancy) and TMR (triple modular redundancy), are also used. Despite these measures, accelerators running DNNs may lack inherent protection.
@@
The vulnerability of DNNs to soft errors, including CNN and transformer-based models, is well-documented [10, 20, 18, 32, 33]. However, previous studies have not extensively explored transformer models in object detection or conducted detailed, large-scale fault injection studies [34, 33]. This study aims to fill this gap by examining the resilience of transformer architectures and exploring effective mitigation strategies against soft errors in these architectures.
@@ 4 Global Clipper
We demonstrate the effectiveness of the Global Clipper solution through two experiments that involve injecting faults into the convolution and linear layers of existing models alongside our proposed method. Using a sample image, we introduce bit flip errors at a random MSB position in the second convolution layer of the ResNet50 backbone within the DINO-DETR model. We then monitor the mean and variance at the 20th ReLU activation layer and the 51st linear layer in the self-attention block immediately following the ResNet50 encoder. This process, illustrated in LABEL:fig:track_mean_var, allows us to evaluate the performance of the Global Clipper and compare it with other mitigation strategies.
We inject faults into the ReLU activation layer in the first experiment and apply various mitigation solutions, as shown in LABEL:fig:track_mean_var_relu. The Ranger method [15] confines data points within the population cluster. In contrast, the Clipper [27], along with the proposed Global Ranger and Global Clipper, more tightly constrain the data to the region of non-faulty inferences. Although all techniques maintain acceptable tolerances, Ranger and Clipper tend to shift data points further from the original non-faulty positions at linear layers.
In our second experiment, we inject faults into the linear layer and activate various mitigation solutions. However, neither Ranger nor Clipper can confine the values to their original positions due to the sensitivity of self-attention layers to faults. Both Ranger and Clipper typically apply restrictions only at the activation layer; however, self-attention includes a SoftMax layer that cannot be similarly restricted without impairing the functionality and accuracy of the block. Hence, introducing the Global Clipper, shown in fig. 3, is essential for protecting attention blocks from faults. With faults introduced into the 51st linear layer, the Global Clipper successfully maintains data points closer to their original positions, as demonstrated in LABEL:fig:track_mean_var_linear.
@@ 5.3 Results: Global Clipper on Transformer models
This section showcases the superior mitigation capability of Global Clipper and Global Hybrid Clipper over existing solutions like Ranger and Clipper. Demonstrated on DINO-DETR and Lite-DETR models with various datasets injected with single bit-flip errors (see LABEL:fig:dino_detr_sdc_ap50 and LABEL:fig:lite_detr_sdc_ap50), we also investigate vulnerability across datasets, including layer-wise fault injection experiments.
The LABEL:fig:dino_detr_sdc_ap50 illustrates the model’s vulnerability and evaluates the effectiveness of the various range restriction techniques in mitigating faults using $IVMOD_{fd}$ and AP50 metrics. This includes results from experiments involving fault injections into weights and neurons. The Global Clipper significantly outperforms the existing solutions like Ranger and Clipper in mitigating the impact of the faults. The Global Clipper performs better in mitigating faults occurring in neurons, reducing faulty detections to nearly 0%, and for weight faults, the vulnerabilities are reduced to less than $\sim$ 3% by outperforming other state-of-the-art algorithms like Ranger and Clipper. Additionally, as shown in LABEL:fig:dino_detr_ap50, a single bit flip in the inferences notably impacts the AP50 metric. For instance, the AP50 of DINO-DETR trained on KITTI decreases from 86.9 to 81.6 when injected with weight faults, and the Global Clipper effectively restores the AP50 to its original accuracy. In the context of DINO-DETR, Global Ranger and Global Clipper demonstrate similar mitigation performance. However, their performance is not as expected when applied to Lite-DETR due to the unique transformer architecture based on DINO-DETR. This presents an interesting scenario, leading to the introduction of the Global Hybrid Clipper, where the Global Clipper is utilized for the Activation layers, and the Global Ranger is employed for the linear layers of self-attention blocks. In this scenario, the hybrid version of Global Clipper restores performance accuracy to baseline and reduces faulty detections significantly, as shown in LABEL:fig:lite_detr_sdc_ap50. For instance, it decreases from 4.5% to 0.5% in the case of neuron faults.

### v3

@@ 1 Motivation
Errors during inference can lead to faulty bounding boxes, significantly altering system behaviour and underscoring the critical need for safer hardware for model execution. DNN accelerators execute models at a high level by constructing a computational graph that uses General matrix-to-matrix multiplication (GEMM) [6] for extensive layer input and weight multiplications. Key components in this process are the Multiply-accumulate (MAC) units within the lower accelerator levels shown in fig. 1 [7]. MAC units in DNN accelerators lack ECC protection, making them particularly vulnerable to soft errors—a major reliability concern. Such errors, often caused by radiation, chip ageing, manufacturing variations, or thermal issues [8, 9, 10], can alter intermediary computational values, leading to incorrect inferences. Research shows that the soft error rate will increase with higher transistor density, reduced feature sizes, and more cores [11, 12, 13]. For example, a 100-core system of 16nm node may fail every 1.5 hours due to soft errors [11], significantly affecting predictions as these propagate through layers, as shown in fig. 1. Although soft errors do not cause permanent damage, they can result in substantial reliability degradation.
@@ 2 Related Work
The reliability of safety-critical DNN models is assessed through various metrics at the application and hardware levels, enhancing their safety and reliability [19, 20, 15, 18]. Research has shown that DNNs are prone to soft errors, with instances of single-bit flips leading to faulty inferences [21, 22, 23, 18]. Traditionally, protection from soft errors in hardware has primarily involved error detection or correction codes (EDC or ECC) [24] for memory and using residuals for computing. These mechanisms are commonly implemented in high-end server-grade CPUs but less so in GPUs due to cost considerations or typical relaxed application requirements. Other techniques based on redundancy, like DMR (dual modular redundancy) and TMR (triple modular redundancy), are also used. Despite these measures, accelerators running DNNs may lack inherent protection. Furthermore, modular redundancy techniques can be implemented through ensembles, as discussed by [25], which may be used to detect and mitigate faults. However, these methods come with substantial computational overhead. To address this issue, budding ensemble solutions [26, 27] could be explored to reduce the computing overhead. Despite their potential, these solutions have not yet been demonstrated for mitigating or detecting soft errors.
@@
The vulnerability of DNNs to soft errors, including a significant number of CNN and few transformer-based models, is well-documented [10, 20, 18, 35, 36, 22]. However, previous studies have not extensively explored transformer models in object detection or conducted detailed, large-scale fault injection studies [37, 36]. This study aims to fill this gap by examining the resilience of transformer architectures and exploring effective mitigation strategies against soft errors in these architectures.
@@ 4 Global Clipper
We demonstrate the effectiveness of the Global Clipper solution using two experiments that involve injecting faults into the convolution and linear layers of existing models alongside our proposed method. Using the sample images from validation datasets, we introduce bit flip errors at a random MSB position in the second convolution layer of the ResNet50 backbone within the DINO-DETR model. We then monitor the mean and variance at the 20th ReLU activation layer and the 51st linear layer in the self-attention block immediately following the ResNet50 encoder. This process, illustrated in LABEL:fig:track_mean_var, allows us to evaluate the performance of the Global Clipper and compare it with other mitigation strategies.
In the first experiment, we inject faults into the ReLU activation layer and apply various mitigation solutions, as shown in LABEL:fig:track_mean_var_relu. The Ranger method [15] confines data points within the extrapolated population cluster space. In contrast, the Clipper [30], along with the proposed Global Clipper, more tightly constrains the data to the region of non-faulty inferences. Although all techniques maintain acceptable tolerances, Ranger and Clipper tend to shift data points further from the original non-faulty positions at linear layers.
In our second experiment, we inject faults into the linear layer and activate various mitigation solutions. However, neither Ranger nor Clipper can confine the values to their original positions due to the sensitivity of self-attention layers to faults. Both Ranger and Clipper typically apply restrictions only at the activation layer; however, self-attention includes a SoftMax layer that cannot be similarly restricted without impairing the functionality and accuracy of the block. Hence, introducing the Global Clipper, shown in fig. 3, is essential for protecting attention blocks from faults. With faults introduced into the 51st linear layer, the Global Clipper successfully maintains data points closer to their original positions, as demonstrated in LABEL:fig:track_mean_var_linear. While the definition of Global Clipper resembles Ranger’s, refining the bounds is crucial as it ensures that feature values remain within the sampled space in specific layers. In both experiments mentioned above, fault injection demonstrated using sample data doesn’t necessarily result in faulty predictions. However, fault locations are sampled to visualize all data points within the area, providing a clear explanation of Global Clipper. Additionally, more faults affect the layer values, causing them to shift into log space compared to the fault-free visualizations. There are a few cases in which Global Clipper may not function out of the box on certain models, requiring slight modifications. These exceptions and adaptations are further elaborated upon in section 5.3.
@@ 5.3 Results: Global Clipper on Transformer models
This section showcases the superior mitigation capability of Global Clipper and Global Hybrid Clipper over existing solutions like Ranger and Clipper. This is demonstrated on DINO-DETR and Lite-DETR models with various datasets injected with single bit-flip errors (see LABEL:fig:dino_detr_sdc_ap50 and LABEL:fig:lite_detr_sdc_ap50), we also investigate vulnerability across datasets, including layer-wise fault injection experiments.
The LABEL:fig:dino_detr_sdc_ap50 illustrates the model’s vulnerability and evaluates the effectiveness of the various range restriction techniques in mitigating faults using $IVMOD_{fd}$ and AP50 metrics. This includes results from experiments involving fault injections into weights and neurons. The Global Clipper significantly outperforms the existing solutions like Ranger and Clipper in mitigating the impact of the faults. The Global Clipper performs better in mitigating faults occurring in neurons, reducing faulty detections to nearly 0%, and for weight faults, the vulnerabilities are reduced to less than $\sim$ 3% by outperforming other state-of-the-art algorithms like Ranger and Clipper. Additionally, as shown in LABEL:fig:dino_detr_ap50, a single bit flip in the inferences notably impacts the AP50 metric. For instance, the AP50 of DINO-DETR trained on KITTI decreases from 86.9 to 81.6 when injected with weight faults, and the Global Clipper effectively restores the AP50 to its original accuracy. In the context of DINO-DETR, Global Ranger and Global Clipper demonstrate similar mitigation performance. However, their performance is not as expected when applied to Lite-DETR due to the unique transformer architecture based on DINO-DETR. This presents an interesting scenario, leading to the introduction of the Global Hybrid Clipper. The Global Hybrid Clipper merges Global Clipper and Ranger layers. In this setup, Global Clipper is applied to Activation layers, while Global Ranger is used for the linear layers within self-attention blocks. In this scenario, the hybrid version of Global Clipper restores performance accuracy to baseline and reduces faulty detections significantly, as shown in LABEL:fig:lite_detr_sdc_ap50. For instance, it decreases from 4.5% to 0.5% in the case of neuron faults.